OpenStudy (anonymous):

Describe how to transform the quantity of the sixth root of x to the fifth power, to the seventh powerinto an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Hero

OpenStudy (campbell_st):

so you have \[(\sqrt[6]{x^5})^7\] is that correct...?

OpenStudy (anonymous):

yessum

OpenStudy (campbell_st):

ok do you know how to write fractional roots for radicals..?

OpenStudy (campbell_st):

\[\sqrt[a]{x^b} = x^{\frac{b}{a}}\]

OpenStudy (campbell_st):

use that law to start then you need the index law for power of a power.... multiply the powers \[(x^a)^b = x^{a \times b}\]

OpenStudy (anonymous):

okay (so x^5/6)^7

OpenStudy (campbell_st):

great so now multiply 5/6 by 7 and that will be the power

OpenStudy (anonymous):

so 35/6?

OpenStudy (anonymous):

@campbell_st